OpenStudy (anonymous):

There are two trucks with width of 3,2m and height 4,1m. They are passing each other in semicircular tunnel with radius 5,5m. How far can the two trucks be from each other without touching tunnel walls? I've been stuck with this since yesterday, thanks for any help! [Answer is 0,94 but i keep getting 0,67 and 0,84..]

OpenStudy (inkyvoyd):

Dude, Europe is awesome.

OpenStudy (dumbcow):

for circle x^2 + y^2 = (5.5)^2 you need to find x value such that y = 4,1 that will tell you how far from the center line the outer side of truck can go without hitting wall --> x^2 = (5.5)^2 - (4.1)^2 --> x^2 = 13.44 --> x = 3.66 subtract width of truck 3.66 - 3.2 = 0.466 this is space each truck can move from center line double it to find how far apart the trucks will be 2*0.466 = 0.932 **sorry i used decimal points, not commas (from USA)

OpenStudy (dumbcow):

weird, not sure how they get 0,94 as answer, even with rounding its 0,93

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (dumbcow):

welcome :)

OpenStudy (anonymous):

i knew the answer but no confidence